# KVQuant: Towards 10 Million Context Length LLM Inference with KV Cache Quantization

### v4

@@ 4.4. Performance Analysis and Memory Savings
5. Conclusion
6. Limitations
7. Acknowledgements.
The authors would like to acknowledge Nicholas Lee for helpful discussions and feedback. We acknowledge gracious support from Intel, Furiosa, Apple, and Samsung SAIT. We also appreciate the support from Microsoft through their Accelerating Foundation Model Research, including great support from Sean Kuno. Furthermore, we appreciate support from Google Cloud, the Google TRC team, and specifically Jonathan Caton, and Prof. David Patterson. Prof. Keutzer’s lab is sponsored by the Intel corporation, Intel One-API, Intel VLAB team, the Intel One-API center of excellence, as well as funding through BDD and BAIR. We appreciate great feedback and support from Ellick Chan, Saurabh Tangri, Andres Rodriguez, and Kittur Ganesh. Sehoon Kim would like to acknowledge the support from the Korea Foundation for Advanced Studies (KFAS). Amir Gholami was supported through funding from Samsung SAIT. Michael W. Mahoney would also like to acknowledge a J. P. Morgan Chase Faculty Research Award as well as the DOE, NSF, and ONR. Our conclusions do not necessarily reflect the position or the policy of our sponsors, and no official endorsement should be inferred.
Appendix A RoPE Equation
@@
(3) $\begin{bmatrix}\mathrm{c}(n\theta_{1})&-\mathrm{s}(n\theta_{1})&\cdots&0&0\\ \mathrm{s}(n\theta_{1})&\mathrm{c}(n\theta_{1})&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\mathrm{c}(n\theta_{d/2})&-\mathrm{s}(n\theta_{d/2})\\ 0&0&\cdots&\mathrm{s}(n\theta_{d/2})&\mathrm{c}(n\theta_{d/2})\\ \end{bmatrix}$
@@
(4) $\begin{bmatrix}x_{1}\\ x_{2}\\ \vdots\\ x_{\frac{d}{2}}\\ x_{\frac{d}{2}+1}\\ \vdots\\ x_{d-1}\\ x_{d}\end{bmatrix}\odot\begin{bmatrix}\mathrm{c}(\theta_{1}n)\\ \mathrm{c}(\theta_{2}n)\\ \vdots\\ \mathrm{c}(\theta_{\frac{d}{2}}n)\\ \mathrm{c}(\theta_{1}n)\\ \vdots\\ \mathrm{c}(\theta_{\frac{d}{2}-1}n)\\ \mathrm{c}(\theta_{\frac{d}{2}}n)\\ \end{bmatrix}+\begin{bmatrix}-x_{\frac{d}{2}+1}\\ -x_{\frac{d}{2}+2}\\ \vdots\\ -x_{d}\\ x_{1}\\ \vdots\\ x_{\frac{d}{2}-1}\\ x_{\frac{d}{2}}\end{bmatrix}\odot\begin{bmatrix}\mathrm{s}(\theta_{1}n)\\ \mathrm{s}(\theta_{2}n)\\ \vdots\\ \mathrm{s}(\theta_{\frac{d}{2}}n)\\ \mathrm{s}(\theta_{1}n)\\ \vdots\\ \mathrm{s}(\theta_{\frac{d}{2}-1}n)\\ \mathrm{s}(\theta_{\frac{d}{2}}n)\\ \end{bmatrix}$
@@
Appendix B Derivation for Sensitivity Analysis
@@
Appendix C Derivation for Quantization Error
@@
(7) $Q(A)^{*}$ $\simeq\operatorname*{arg\,min}_{Q}\sum_{i=1}^{N}\mathcal{F}_{ii}\big{(}A_{i}-Q(A_{i})\big{)}^{2}$
(8) $=\operatorname*{arg\,min}_{Q}\sum_{i=1}^{N}\mathcal{F}_{ii}\Big{(}s_{i}^{2}\big{(}A_{i,norm}-Q(A_{i,norm})\big{)}^{2}\Big{)}$
Appendix D Key and Value Dynamic Range
@@
Figure 5. Distribution of the magnitude of elements of Key (Pre-RoPE) and Value activations for different layers of LLaMA-7B, computed on a single sample with sequence length 2K from the Wikitext-2 dataset. The normalized magnitude is computed by dividing by the largest magnitude value in that layer. As one can see, for both Key and Value activations, the majority of values lie in a small portion of the dynamic range, with a few numerical outliers skewing the dynamic range (and thereby reducing the fidelity when quantizing to low precision).
Appendix E Per-Channel Key Quantization Ablations
@@
Table 7. Ablation Study: Perplexity comparison of per-token and per-channel quantization for KV cache activations for LLaMA-7B. PT refers to per-token quantization, and PC refers to per-channel quantization.
@@
Appendix F Pre-RoPE Key Quantization Ablations
@@
Table 8. Ablation Study: Perplexity comparison of Pre-RoPE and post-RoPE Key quantization for LLaMA-7B (using per-channel Key quantization and per-token Value quantization). Pre-RoPE quantization leads to significant improvement (see Section 3.2 for more details).
@@
Appendix G Sensitivity-Weighted Non-Uniform Quantization Ablations
@@
Table 9. Ablation Study: Ablation of our sensitivity-weighted non-uniform datatype for different models on Wikitext-2. All experiments use pre-RoPE per-channel quantization for Keys and per-token quantization for Values (meaning that all configurations are the same as in KVQuant, except for the datatype). We compare against both uniform (int3) and non-uniform (nf3) [8] approaches, as well as with using “unweighted” k-means (i.e., not sensitivity-weighted) and “Fisher-weighted k-means” (without accounting for per-channel scaling factors) to compute the non-uniform quantization signposts. Note that there is slight variation in average bitwidth across models due to the differing hidden dimensions. Results report perplexity with 2K/4K/8K sequence length for LLaMA, Llama-2, and Mistral, respectively.
@@
Appendix H Per-Vector Dense-and-Sparse Quantization Ablations
@@
Table 10. Ablation Study: Perplexity comparison of different outlier isolation methods for LLaMA-7B on Wikitext-2. Per-vector outlier detection allows for significant accuracy improvements relative to per-tensor outlier detection. All experiments use per-token quantization for Values and per-channel quantization for Keys (pre-RoPE). “PV” refers to using per-vector outlier thresholds, and “PM” refers to using a single per-matrix outlier threshold.
@@
Appendix I Attention Sink-Aware Quantization Ablations
@@
Table 11. Ablation Study: Perplexity with and without Attention Sink-Aware quantization on Wikitext-2 with various models. Attention Sink-Aware quantization consistently improves perplexity across different models and bit widths, particularly with lower bit widths and without sparsity.
@@
Appendix J Calibration Ablations
@@
Table 12. Ablation Study: Model accuracy when using offline calibration for Keys with LLaMA-7B. When incorporating outlier detection, offline calibration for Keys is able to perform comparably with online calibration. All nuq3 experiments use per-token quantization for Values and per-channel quantization for Keys (pre-RoPE), and experiments with outliers use per-vector outlier detection.
@@
Table 13. $topk$ runtime on a vector of length 4096 for computing outlier thresholds when using 1% sparsity (compared with the runtime for the QKV matrix multiplications, which are 4096$\times$4096 by 4096 matrix-vector multiplications for the LLaMA-7B model). The runtime is reported on a system with an A6000 GPU and an Intel Xeon Gold 6126 CPU. We find that the runtime for the $topk$ operation is only 45% of the runtime of each matvec operation. Additionally, the $topk$ operation can be performed efficiently on the CPU; we can therefore run this operation in parallel with subsequent linear layer operations on the GPU to compress the activations dynamically without added overhead.
@@
Table 14. Runtime for computing Fisher information as well as for calibration (including k-means) with 16 samples for LLaMA-65B quantization. Runtime for computing Fisher information was computed on an 8-GPU A100-80GB system. Runtime for calibration (including k-means) was performed on an Intel Xeon Gold 6442Y CPU, and is shown for a single layer. Note that calibration is independent for each layer, so it can be easily parallelized.
@@
Appendix K Additional Experimental Details
For our empirical evaluation, we use 16 calibration samples of sequence length 2K from the Wikitext-2 training set (as well as the corresponding gradients) to derive the per-channel scaling factors and zero-points, and to derive the non-uniform datatypes for both Keys and Values. We measured perplexity on both Wikitext-2 and on C4 using a sequence length equal to the maximum context length of the model (2K for LLaMA, 4K for Llama-2, and 8K for Llama-3 and Mistral-7B). For baseline experiments, we use post-RoPE quantization, both since this is required from an efficiency perspective without a dedicated kernel implementation, and because it provides better accuracy when quantizing Keys per-token as shown in Appendix N.
@@
Appendix L Comparison Between Different Datatypes and Sparsity Levels
@@
Table 15. Comparison of our NUQ datatype with and without sparsity against other data types on different models using the perplexity (PPL) measured on Wikitext-2. Non-uniform quantization (“nuq”) results are using pre-RoPE per-channel quantization for Keys. “gs64/128” refers to baseline experiments using grouping with group size 64/128. Note that there is a slight variation in average bitwidth across models due to the differing hidden dimensions.
@@
Appendix M Full Perplexity Evaluation
Tables 16 and 17 show perplexity evaluation across all LLaMA, Llama-2, Llama-3, and Mistral models on Wikitext-2 and C4, respectively. These results demonstrate the benefits of our approach for KV cache compression across different models and bit widths.
Table 16. Full evaluation of our method for different models on all LLaMA, Llama-2, Llama-3, and Mistral models using the perplexity on Wikitext-2. KVQuant results are using pre-RoPE per-channel quantization for Keys. Average bit widths assume a sequence length of 128K (ignoring context length limits for the models). Note that ATOM and FlexGen use 4-bit quantization with group sizes of 128 and 64 with uniform quantization, respectively, and we extend their methods to 3-bit and 2-bit quantization. We leverage Attention-Sink Aware quantization for all bit widths. We used post-RoPE quantization for all baseline methods since it achieves higher accuracy when quantizing Keys per-token as shown in Appendix N.
@@
Table 17. Full evaluation of our method for different models on all LLaMA, Llama-2, Llama-3, and Mistral models using the perplexity on C4. KVQuant results are using pre-RoPE per-channel quantization for Keys. Average bit widths assume a sequence length of 128K (ignoring context length limits for the models). Note that ATOM and FlexGen use 4-bit quantization with group sizes of 128 and 64 with uniform quantization, respectively, and we extend their methods to 3-bit and 2-bit quantization. We leverage Attention-Sink Aware quantization for all bit widths. We used post-RoPE quantization for all baseline methods since it achieves higher accuracy when quantizing Keys per-token as shown in Appendix N.
@@
Appendix N Post-RoPE Per-Token Quantization Ablation
@@
Table 18. Model accuracy when using pre-RoPE and post-RoPE quantization for LLaMA-7B with per-token Key quantization. Our experiments demonstrate that post-RoPE quantization is superior when using per-token Key quantization. Therefore, we decided to use these results for baseline comparison with per-token quantization.
@@
Appendix O Additional Long Sequence Length Evaluation
Table 19 shows evaluation on Longbench [3] for the LLaMA-2-7B-32K model. The max input context length is set at 31500, and results using KIVI are also included for reference [25]. Note that we only ran LongBench evaluation using the 32K model since the majority of datasets have an average length greater than the context length of the non-extended models (i.e., 4K). The average input length was 12.2K tokens when using the LLaMA tokenizer (and truncating to fit within 32K context length), and hence most inputs would be heavily truncated for shorter context length models. Our results demonstrate that our method has little degradation relative to baseline fp16 inference, as well as improved accuracy for similar average bit width relative to KIVI.
Table 19. LongBench evaluation for the LLaMA-2-7B-32K model using KVQuant-3bit-1%. Comparisons with KIVI are included for reference, using the configuration with group size of 32 and 128-element fp16 residual [25]. Average bit widths are estimated for each approach assuming 12.2K context length, which was the average number of tokens across all tasks.
Task fp16 Baseline KIVI-2-gs32-r128 KVQuant-3bit-1%
Average Bit Width 16 3.17 3.33
Qasper 10.51 10.66 13.67
QMSum 16.99 16.36 16.53
MultiNews 22.15 23.37 16.54
TREC 71 71 71
TriviaQA 87.79 80.8 87.55
SAMSum 43.97 43.93 43.95
LCC 62.14 60.61 61.52
RepoBench-P 59.99 57.74 59.5
NarrativeQA 17.96 19.25 18.87
2WikiMQA 12.53 11.19 12.55
Musique 6.19 6.38 6.25
HotPotQA 12.55 12.48 12.07
Passage Retrieval 23 13.58 19.5
MultifieldQA-en 33.43 24.78 30.93
GovReport 29.65 27.05 27.1
Passage Count 1.50 1.50 1.75
Average Score 31.96 30.04 31.21
Appendix P Kernel Implementation Details
@@
Table 20. Average latency (in microseconds) for the Key and Value dense nuq4 kernels as well as for the sparse kernels (with 1% outliers), benchmarked on an A6000 GPU for the LLaMA-2-7B-32K model. Benchmarking results are reported for different sequence lengths ($l$). fp16 matrix-vector multiplication latencies are included for reference, and the Key multiplication time also includes the time to apply RoPE to the newly appended Key vector. We find that our dense-and-sparse approach (even with 1% outliers) provides latency benefits relative to the fp16 baseline, even when accounting for the time to compress activations online.

### v5

@@ 16 $\rightarrow$ 2-bit 32K 128K 1M 10M (16 $\rightarrow$ 2-bit)
Appendix B Additional Related Works
B.1 Outlier-Aware LLM Quantization
LLMs have been known to have distinct outliers both in weights and activations [7, 9, 17]. SqueezeLLM and SpQR both decompose the weight matrix into a sparse matrix containing a small portion of outliers and a dense matrix that can be accurately quantized to low precision (referred to as dense-and-sparse or sparse-quantized representation) [9, 17]. LLM.int8() [7] handled particular outlier channels separately in higher precision, and SmoothQuant [41] migrates quantization difficulty due to outlier channels to weights in order to support joint weight-activation quantization. Other works reconsidered the dimension along which we quantize in order to reduce quantization error (or else added per-channel compensation to improve quantization performance) [4, 14, 39, 38]. In this work, we demonstrate that per-channel pre-RoPE Key quantization provides significant accuracy benefits given the outlier structure in Keys, and that dense-and-sparse quantization can be efficiently applied for KV cache quantization.
B.2 Non-uniform LLM Quantization
Non-uniform quantization has also been applied in the context of LLMs. Non-uniform quantization allows for more flexible quantization signpost placement relative to uniform quantization methods, enabling improved accuracy for the same bit precision [17, 8]. Building on the observation that model parameters tend to be approximately normally-distributed, prior work has proposed the NormalFloat datatype [8]. SqueezeLLM [17] derived per-channel non-uniform quantization signposts using a sensitivity-weighted k-means approach. In this work, we show that we can derive accurate per-layer non-uniform datatypes using a sensitivity-weighted k-means approach with KV cache activations.
Appendix C RoPE Equation
@@
$\begin{bmatrix}\mathrm{c}(n\theta_{1})&-\mathrm{s}(n\theta_{1})&\cdots&0&0\\ \mathrm{s}(n\theta_{1})&\mathrm{c}(n\theta_{1})&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\mathrm{c}(n\theta_{d/2})&-\mathrm{s}(n\theta_{d/2})\\ 0&0&\cdots&\mathrm{s}(n\theta_{d/2})&\mathrm{c}(n\theta_{d/2})\\ \end{bmatrix}$ (2)
@@
$\begin{bmatrix}x_{1}\\ x_{2}\\ \vdots\\ x_{\frac{d}{2}}\\ x_{\frac{d}{2}+1}\\ \vdots\\ x_{d-1}\\ x_{d}\end{bmatrix}\odot\begin{bmatrix}\mathrm{c}(\theta_{1}n)\\ \mathrm{c}(\theta_{2}n)\\ \vdots\\ \mathrm{c}(\theta_{\frac{d}{2}}n)\\ \mathrm{c}(\theta_{1}n)\\ \vdots\\ \mathrm{c}(\theta_{\frac{d}{2}-1}n)\\ \mathrm{c}(\theta_{\frac{d}{2}}n)\\ \end{bmatrix}+\begin{bmatrix}-x_{\frac{d}{2}+1}\\ -x_{\frac{d}{2}+2}\\ \vdots\\ -x_{d}\\ x_{1}\\ \vdots\\ x_{\frac{d}{2}-1}\\ x_{\frac{d}{2}}\end{bmatrix}\odot\begin{bmatrix}\mathrm{s}(\theta_{1}n)\\ \mathrm{s}(\theta_{2}n)\\ \vdots\\ \mathrm{s}(\theta_{\frac{d}{2}}n)\\ \mathrm{s}(\theta_{1}n)\\ \vdots\\ \mathrm{s}(\theta_{\frac{d}{2}-1}n)\\ \mathrm{s}(\theta_{\frac{d}{2}}n)\\ \end{bmatrix}$ (3)
@@
Appendix D Derivation for Sensitivity Analysis
@@
Appendix E Derivation for Quantization Error
@@
$Q(A)^{*}$ $\simeq\operatorname*{arg\,min}_{Q}\sum_{i=1}^{N}\mathcal{F}_{ii}\big{(}A_{i}-Q(A_{i})\big{)}^{2}$ (6)
$=\operatorname*{arg\,min}_{Q}\sum_{i=1}^{N}\mathcal{F}_{ii}\Big{(}s_{i}^{2}\big{(}A_{i,norm}-Q(A_{i,norm})\big{)}^{2}\Big{)}$ (7)
Appendix F Key and Value Dynamic Range
@@
Figure 4: Distribution of the magnitude of elements of Key (Pre-RoPE) and Value activations for different layers of LLaMA-7B, computed on a single sample with sequence length 2K from the Wikitext-2 dataset. The normalized magnitude is computed by dividing by the largest magnitude value in that layer. As one can see, for both Key and Value activations, the majority of values lie in a small portion of the dynamic range, with a few numerical outliers skewing the dynamic range (and thereby reducing the fidelity when quantizing to low precision).
Appendix G Per-Channel Key Quantization Ablations
@@
Table 9: Ablation Study: Perplexity comparison of per-token and per-channel quantization for KV cache activations for LLaMA-7B. PT refers to per-token quantization, and PC refers to per-channel quantization.
@@
Appendix H Pre-RoPE Key Quantization Ablations
@@
Table 10: Ablation Study: Perplexity comparison of Pre-RoPE and post-RoPE Key quantization for LLaMA-7B (using per-channel Key quantization and per-token Value quantization). Pre-RoPE quantization leads to significant improvement (see Section 3.2 for more details).
@@
Appendix I Sensitivity-Weighted Non-Uniform Quantization Ablations
@@
Table 11: Ablation Study: Ablation of our sensitivity-weighted non-uniform datatype for different models on Wikitext-2. All experiments use pre-RoPE per-channel quantization for Keys and per-token quantization for Values (meaning that all configurations are the same as in KVQuant, except for the datatype). We compare against both uniform (int3) and non-uniform (nf3) [8] approaches, as well as with using “unweighted” k-means (i.e., not sensitivity-weighted) and “Fisher-weighted k-means” (without accounting for per-channel scaling factors) to compute the non-uniform quantization signposts. Note that there is slight variation in average bitwidth across models due to the differing hidden dimensions. Results report perplexity with 2K/4K/8K sequence length for LLaMA, Llama-2, and Mistral, respectively.
@@
Appendix J Per-Vector Dense-and-Sparse Quantization Ablations
@@
Table 12: Ablation Study: Perplexity comparison of different outlier isolation methods for LLaMA-7B on Wikitext-2. Per-vector outlier detection allows for significant accuracy improvements relative to per-tensor outlier detection. All experiments use per-token quantization for Values and per-channel quantization for Keys (pre-RoPE). “PV” refers to using per-vector outlier thresholds, and “PM” refers to using a single per-matrix outlier threshold.
@@
Appendix K Attention Sink-Aware Quantization Ablations
@@
Table 13: Ablation Study: Perplexity with and without Attention Sink-Aware quantization on Wikitext-2 with various models. Attention Sink-Aware quantization consistently improves perplexity across different models and bit widths, particularly with lower bit widths and without sparsity.
@@
Appendix L Calibration Ablations
Figure 5 outlines the accuracy and efficiency challenges which our work addresses in order to enable accurate and efficient KV cache quantization. We demonstrate that we can circumvent the challenges related to efficiently performing online per-channel scaling factor computation by instead calibrating offline for the scaling factor without hurting accuracy. Additionally, we show that we can perform per-token scaling factor and outlier threshold computation efficiently online during inference, thereby enabling accurate per-token quantization without compromising on efficiency.
Figure 5: One typically achieves better performance when the scaling factor/zero point are computed online. However, this is quite challenging to do for per-channel quantization, as these factors will not only need to be recomputed for every new Key appended to the Key cache, but also all the prior cached Keys will need to be updated. As such, we use a calibration set to compute per-channel scaling factors offline. A similar challenge exists for per-token quantization, but online calibration for this does not require updating prior cached Values. In Section 3.6 and Appendix L, we discuss how we are able to efficiently compute outlier thresholds / scaling factors for per-token calibration, thereby enabling online computation.
@@
Table 14: Ablation Study: Model accuracy when using offline calibration for Keys with LLaMA-7B. When incorporating outlier detection, offline calibration for Keys is able to perform comparably with online calibration. All nuq3 experiments use per-token quantization for Values and per-channel quantization for Keys (pre-RoPE), and experiments with outliers use per-vector outlier detection.
@@
Table 15: $topk$ runtime on a vector of length 4096 for computing outlier thresholds when using 1% sparsity (compared with the runtime for the QKV matrix multiplications, which are 4096$\times$4096 by 4096 matrix-vector multiplications for the LLaMA-7B model). The runtime is reported on a system with an A6000 GPU and an Intel Xeon Gold 6126 CPU. We find that the runtime for the $topk$ operation is only 45% of the runtime of each matvec operation. Additionally, the $topk$ operation can be performed efficiently on the CPU; we can therefore run this operation in parallel with subsequent linear layer operations on the GPU to compress the activations dynamically without added overhead.
@@
Table 16: Runtime for computing Fisher information as well as for calibration (including k-means) with 16 samples for LLaMA-65B quantization. Runtime for computing Fisher information was computed on an 8-GPU A100-80GB system. Runtime for calibration (including k-means) was performed on an Intel Xeon Gold 6442Y CPU, and is shown for a single layer. Note that calibration is independent for each layer, so it can be easily parallelized.
@@
Appendix M Additional Experimental Details
For our empirical evaluation, we use 16 calibration samples of sequence length 2K from the Wikitext-2 training set (as well as the corresponding gradients) to derive the per-channel scaling factors and zero-points, and to derive the non-uniform datatypes for both Keys and Values. While we use KVQuant models that are calibrated on the Wikitext-2 dataset for all experiments, in Appendix L, we include an additional experiment that demonstrates the robustness of the calibration process to the choice of data.
We measured perplexity on both Wikitext-2 and on C4 using a sequence length equal to the maximum context length of the model (2K for LLaMA, 4K for Llama-2, and 8K for Llama-3 and Mistral-7B). For generative tasks, when processing the input prompt, the Key/Value matrix multiplications are computed using the fp16 Keys and Values, and then they are separately compressed into low precision. For baseline experiments, we use post-RoPE quantization, both since this is required from an efficiency perspective without a dedicated kernel implementation, and because it provides better accuracy when quantizing Keys per-token as shown in Appendix P.
@@
Appendix N Comparison Between Different Datatypes and Sparsity Levels
@@
Table 17: Comparison of our NUQ datatype with and without sparsity against other data types on different models using the perplexity (PPL) measured on Wikitext-2. Non-uniform quantization (“nuq”) results are using pre-RoPE per-channel quantization for Keys. “gs64/128” refers to baseline experiments using grouping with group size 64/128. Note that there is a slight variation in average bitwidth across models due to the differing hidden dimensions.
@@
Appendix O Full Perplexity Evaluation
Tables 18 and 19 show perplexity evaluation across all LLaMA, Llama-2, Llama-3, and Mistral models on Wikitext-2 and C4, respectively. These results demonstrate the benefits of our approach for KV cache compression across different models and bit widths.
Table 18: Full evaluation of our method for different models on all LLaMA, Llama-2, Llama-3, and Mistral models using the perplexity on Wikitext-2. KVQuant results are using pre-RoPE per-channel quantization for Keys. Average bit widths assume a sequence length of 128K (ignoring context length limits for the models). Note that ATOM and FlexGen use 4-bit quantization with group sizes of 128 and 64 with uniform quantization, respectively, and we extend their methods to 3-bit and 2-bit quantization. We leverage Attention-Sink Aware quantization for all bit widths. We used post-RoPE quantization for all baseline methods since it achieves higher accuracy when quantizing Keys per-token as shown in Appendix P.
@@
Table 19: Full evaluation of our method for different models on all LLaMA, Llama-2, Llama-3, and Mistral models using the perplexity on C4. KVQuant results are using pre-RoPE per-channel quantization for Keys. Average bit widths assume a sequence length of 128K (ignoring context length limits for the models). Note that ATOM and FlexGen use 4-bit quantization with group sizes of 128 and 64 with uniform quantization, respectively, and we extend their methods to 3-bit and 2-bit quantization. We leverage Attention-Sink Aware quantization for all bit widths. We used post-RoPE quantization for all baseline methods since it achieves higher accuracy when quantizing Keys per-token as shown in Appendix P.
@@
Appendix P Post-RoPE Per-Token Quantization Ablation
@@
Table 20: Model accuracy when using pre-RoPE and post-RoPE quantization for LLaMA-7B with per-token Key quantization. Our experiments demonstrate that post-RoPE quantization is superior when using per-token Key quantization. Therefore, we decided to use these results for baseline comparison with per-token quantization.
@@
Appendix Q Experiments on Calibration Data Robustness
To evaluate the robustness of our quantization method to the choice of calibration datasets, we measure the perplexity on Wikitext-2 and C4 using different quantized models calibrated with Wikitext-2 and C4. As shown in Table 21, the resulting perplexity numbers remain similar even when the models are calibrated with out-of-domain examples (e.g., calibrated with Wikitext-2 and evaluated on C4, and vice versa), demonstrating the robustness of the calibration process to the choice of data.
Table 21: Perplexity (PPL) results on Wikitext-2 and C4 using different quantization schemes, calibrated using Wikitext-2 and C4.
Datatype Wikitext-2 C4
Calib. with Wikitext-2 Calib. with C4 Calib. with Wikitext-2 Calib. with C4
4-bit, 1% sparsity 5.69 5.70 7.09 7.09
3-bit, 1% sparsity 5.75 5.75 7.13 7.13
2-bit, 1% sparsity 6.05 6.07 7.38 7.38
Appendix R Kernel Implementation Details
@@
Table 22: Average latency (in microseconds) for the Key and Value dense nuq4 kernels as well as for the sparse kernels (with 1% outliers), benchmarked on an A6000 GPU for the LLaMA-2-7B-32K model. Benchmarking results are reported for different sequence lengths ($l$). fp16 matrix-vector multiplication latencies are included for reference, and the Key multiplication time also includes the time to apply RoPE to the newly appended Key vector. We find that our dense-and-sparse approach (even with 1% outliers) provides latency benefits relative to the fp16 baseline, even when accounting for the time to compress activations online.